6. The Agnes Co. is considering the acquisition of the Bilko Co. Agnes projects that earnings next year will be $500. Agnes has 1000 shareholders and a P/E ratio of 10. Bilko projects that earnings next year will be $200. Bilko has 2000 shareholders and a P/E ratio of 15. If the acquisition is completed the new Agnes-Bilko (AB) Co. earnings are expected to be $800 (Note: the earnings of AB is not the sum of A and B).

a. At what exchange ratio (ER) will the old Agnes shareholders suffer 15% dilution in EPS?

b. At what ER will the old Bilko shareholders gain 10% in EPS?

c. If the actual ER is .30 and the new AB has a P/E ratio of 15, what will be the total value of the new AB?

I need help with Parts a, b and c please.

Thank you
Answer #1

Ans:

a) EPS = Earrings/ no of shares

= 500/1000

= 0.5

15 % dilution in EPS = 0.5*85/100

= 0.425

P/E Ratio = price per share / EPS

10 = price per share / 0.425

Price per share = 10 / 0.425

= $ 23.53

Exchange Ratio = Market share price of B / Market share price of A

Share price of B = P/E ratio / EPS

EPS = 200 / 2000 = 0.1

Share price of B = 15 / 0.1

= $ 150

Exchange Ratio of Agnes = 150 / 23.53

= 6.37

b) 10% gain in EPS of Bilko = 0.1 * 110/100

= 0.11

Share price of B = 15 / 0.11 = 136.36

Exchange Ratio = market share price of A / market share price of B

Share price of A = P/E Ratio / EPS

= 10 / 0.5 = $ 20

Exchange Ratio of Bilko = 20 / 136.36

. = 0.15 ( rounded off)

c) P/E Ratio = price per share / EPS

EPS = 800 / 1000+2000

= 0.267

15 = price per share / 0.267

Price per share = 15 / 0.267

= 56.17

Total value of AB = Price per share * no of shares

= 56.17 * 3000

= $ 168510